Mike and Colin run around a track a 450 meter track in opposite directions. They start at the same point and start at the same time. Mike's speed is 2.0 m/s and Colins speed is 2.5 m/s. How far will colin run when they meet?
Answer by checkley77(12844)

450=2t+2.5t
450=4.5t
t=450/4.5
t=100 running time
2.5*t=250 meters for Colin.
2*100=200 meters for Mike.